OpenStudy (anonymous):

|k+9|=4 ? give answer for a medal and ill be a fan

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

-5+9=?

OpenStudy (ybarrap):

$$ |k+9|=4\\ (k+9)^2=4^2=16\\ k^2+18k+81=16\\ k^2+18k+81-16=0\\ k^2+18k+65\\ (k+5) (k+13) = 0 $$ So k=-5 and k=-13. Make sense?